OpenStudy (anonymous):

obtain 5+2i / 3-i in the form of a+ib , a, b belongto R

OpenStudy (anonymous):

multiply by (3+i)/((3+i) & simplify

OpenStudy (anonymous):

\[(5+2i)(3+i)/(3-i)(3+i)=(15+5i+6i-2)/(9+1)=(13+11i)/10=

OpenStudy (anonymous):

=13/10 +11/10 *i

OpenStudy (anonymous):

are you good?

OpenStudy (anonymous):

ok thanks

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

please check:'Good Answer' button" to remove this question from unanswered :)